40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Aug 5 11:04:07 2017, gautam, Update, SUS, MC1 glitches return MC1_glitch_Aug42017.pngMC1_glitch_zoomed.png
    Reply  Wed Aug 9 15:15:47 2017, gautam, Update, SUS, MC1 glitches return MC1_glitch.png
       Reply  Thu Aug 10 21:01:43 2017, gautam, Update, SUS, MC1 glitches debugging MC_QUAD_10AUG2017.jpgMCR_10AUG2017.jpg
          Reply  Fri Aug 11 12:32:46 2017, gautam, Update, SUS, MC1 glitches debugging MC1_freeswinging.pdfMC1_inmatrix.pngMC1_sensors.png
             Reply  Fri Aug 11 17:36:47 2017, gautam, Update, SUS, MC1 <--> MC3 MC1_glitch_watchdog_shutdown.png
                Reply  Sat Aug 12 14:09:36 2017, gautam, Update, SUS, Glitches stay on MC1 MC1_glitch.pngMC_12hr_trend.pngMC1_glitches_intermittent.png
                   Reply  Sat Aug 12 22:35:22 2017, rana, Update, SUS, Glitches stay on MC1 
                      Reply  Sun Aug 13 01:35:09 2017, Koji, Update, SUS, Glitches stay on MC1 
                         Reply  Mon Aug 14 20:01:38 2017, gautam, Update, SUS, Glitches stay on MC1 
                Reply  Wed Aug 16 19:50:17 2017, gautam, Update, SUS, MC1 <--> MC3 switched back 
                   Reply  Thu Aug 17 11:17:49 2017, gautam, Update, SUS, MC1 <--> MC3 switched back MC1_misaligned.pngMC1_glitch.png
                      Reply  Thu Aug 17 17:33:01 2017, gautam, Update, SUS, MC1 <--> MC3 switched back 
Message ID: 13168     Entry time: Sat Aug 5 11:04:07 2017     Reply to this: 13178
Author: gautam 
Type: Update 
Category: SUS 
Subject: MC1 glitches return 

See Attachment #1, which is full (2048Hz) data for a 3 minute stretch around when I saw the MC1 glitch. At the time of the glitch, WFS loops were disabled, so the only actuation on MC1 was via the local damping loops. The oscillations in the MC2 channels are the autolocker turning on the MC2 length tickle.

Nikhil and I tried the usual techniques of squishing cables at the satellite box, and also at 1X4/1X5, but the glitching persists. I will try and localize the problem this weekend. This thread details investigations the last time something like this happened. In the past, I was able to fix this kind of glitching by replacing the (high speed) current buffer IC LM6321M. These are present in a two places: Satellite box (for the shadow sensor LED current drive), and on the coil driver boards. I think we can rule out the slow machine ADCs that supply the static PIT and YAW bias voltages to the optic, as that path is low-passed with a 4th order filter @1Hz, while the glitches that show up in the OSEM sensor channels do not appear to be low-passed, as seen in the zoomed in view of the glitch in Attachment #2 (but there is an LM6321 in this path as well).

Attachment 1: MC1_glitch_Aug42017.png  55 kB  | Hide | Hide all
MC1_glitch_Aug42017.png
Attachment 2: MC1_glitch_zoomed.png  32 kB  Uploaded Sun Aug 6 16:07:06 2017  | Hide | Hide all
MC1_glitch_zoomed.png
ELOG V3.1.3-